Five years ago, my age was half of the age I will be in 8 years. How old I am now?
Answer by ankor@dixie-net.com(22740)

Five years ago, my age was half of the age I will be in 8 years. How old I am now?
:
Let x = present age
then
(x-5) = age 5 yrs ago
and
(x+8) = age in 8 yrs
:
write an equation for the statement:
"Five years ago, my age was half of the age I will be in 8 years."
(x-5) = .5(x+8)
x - 5 = .5x + 4
x -.5x = 4 + 5
.5x = 9
x = 9/.5
x = 18 yrs is the present age
:
Check solution in the above statement:
13 = .5*26